The editor of the school newspaper must reduce the size of a graph to fit in one column.the original graph is 2 inches by 2 inches, and the scale factor from the original to the reduced graph is 8:3.determine the dimensions of the graph as it will appear in one column of the newspaper.
Answer by stanbon(75887)

The editor of the school newspaper must reduce the size of a graph to fit in one column.the original graph is 2 inches by 2 inches, and the scale factor from the original to the reduced graph is 8:3.determine the dimensions of the graph as it will appear in one column of the newspaper.
-----------
Solve x/2 = 3/8
x = 6/8 = 3/4
Ans: 2 by 2(3/8) = 2 by 6/8 = 2" by (3/4)"
------
Cheers,
Stan H.
--------------